# A Vision-Language Pre-training Model-Guided Approach for Mitigating Backdoor Attacks in Federated Learning

### v1

@@ Conclusion
In this paper, we proposed CLIP-Fed that is a general backdoor defense scheme guided by the VLP model CLIP. Without accessing client data or assuming homogeneous distributions, CLIP-Fed enhanced effectiveness by combining pre-aggregation malicious model filtering with post-aggregation purification. Leveraging MLLMs and frequency analysis, we constructed and enhanced a server dataset, addressing privacy concerns and improving coverage of diverse backdoor triggers with minimal utility loss. We further proposed a global model purification strategy based on contrastive learning and knowledge distillation, which corrected class prototype deviations and weakened the link between triggers and target labels, thus improving robustness against trigger samples. Extensive experiments on multiple datasets demonstrated that CLIP-Fed outperformed existing methods in defending against backdoor attacks under varying data distributions and malicious proportions of clients.

### v2

@@ 5. Conclusion
In this paper, we proposed CLIP-Fed that is a general novel backdoor defense scheme guided by the VLP model CLIP. Without accessing client data or assuming homogeneous distributions, CLIP-Fed enhanced overall effectiveness by combining pre-aggregation malicious model filtering with post-aggregation purification. Leveraging MLLMs and frequency analysis, we constructed and enhanced a server dataset, addressing privacy concerns and improving coverage of diverse backdoor triggers with minimal utility loss. We further proposed a global model purification strategy based on contrastive learning and knowledge distillation, which corrected class prototype deviations and weakened the link between triggers and target labels, thus improving robustness against potential trigger samples. Extensive experiments on multiple datasets demonstrated that CLIP-Fed outperformed existing methods in defending against backdoor attacks under significantly varying data distributions and malicious proportions of clients.